Can you put these into y=mx+b
5x+y=-4 and x-y=2
-0.45-y=1.35 and -1.8+y=-1.8
-0.4x=0.8y=-16 and 1.2x+0.4y=1

Answer by stanbon(75887)

put these into y=mx+b
5x+y=-4 and x-y=2
y = -5x-4 and y = x-2
----------------------------------
-0.45-y=1.35 and -1.8+y=-1.8
y = -1.8 and y = 0
----------------------------------
-0.4x=0.8y=-16 and 1.2x+0.4y=1
Too many = signs and y = (-1/2)x + (5/2)
==========================================
Cheers,
Stan H.
=========================